OpenStudy (anonymous):

Help on these two problems please!! 1. tan^-1 (-0.09) 2. cos theta- tan theta(cos theta)=0

OpenStudy (anonymous):

@jim_thompson5910 ?

OpenStudy (anonymous):

the first one is a calculator exercise

OpenStudy (anonymous):

when i plug it in i get -5.14

OpenStudy (anonymous):

i dont know what to do with that number though

OpenStudy (anonymous):

the second one factors as \[\cos(\theta)(1-\tan(\theta))=0\] solve for each factor \[\cos(\theta)=0\] and \[1-\tan(\theta)=0\]

OpenStudy (anonymous):

don't do anything with it, that is your answer

OpenStudy (anonymous):

thanks!